OpenStudy (anonymous):

You want to put a fence along the front edge of your yard (length). You know that the length (L) is 10 feet longer than the width (W), and the perimeter is 85 feet. How long is the front edge of your yard? Set-up the following systems of the equation, don't need solve.

OpenStudy (binary3i):

let the length be L and width be W so L=W+10 W=L-10 L+W=85 L+ L -10=85 2L=85+10 L=95/2=47.5

OpenStudy (binary3i):

L+W=perimeter and L=W+10

OpenStudy (binary3i):

yeah that will give you W using the value of W you can get L

OpenStudy (binary3i):

yes

OpenStudy (binary3i):

you'r welcome